OpenStudy (anonymous):

Condense to a single logarithm. logv3x+logv3y+4logv3^z

OpenStudy (superdavesuper):

u can use log a + log b = log (a*b) to simplify

OpenStudy (superdavesuper):

oh also a*(log b) = log (b^a)

OpenStudy (anonymous):

what would that look like? I'm not sure where the numbers plug into the equation

OpenStudy (superdavesuper):

u got logv3x+logv3y and i show log a + log b = log (a*b) so a=x and b=3y...

OpenStudy (superdavesuper):

ooops typo i mean a=3x n b=3y